OpenStudy (rebeccapink6398):

@31356 :D Use a ratio to compare the values. Simplify the ratio. 30 in. to 2 ft A. 15 : 1 B. 5 : 4 C. 4 : 5 D. 1 : 15

OpenStudy (31356):

Thanks for the medal! :D

OpenStudy (cwrw238):

first convert 2 ft to inches

OpenStudy (31356):

Just simplify and you will find the answer :D

OpenStudy (rebeccapink6398):

lol Thx! :d how did u get that?

OpenStudy (rebeccapink6398):

OH! ok

OpenStudy (31356):

Oh yea, wait convert first

OpenStudy (anonymous):

Nice explanation (not) 30 in. : 2 ft 30 in : 24 in ........... 2 x 12in 5 : 4 divide both sides by 6

OpenStudy (31356):

Yea, didn't read the question carefully, lol

OpenStudy (31356):

Which then is B) 5/4

OpenStudy (31356):

Sorry about that o.0

OpenStudy (rebeccapink6398):

lol its ok

OpenStudy (31356):

Should of read the question carefully first, lol o.0

OpenStudy (31356):

Got to go now, I will answer your other answer later?

OpenStudy (rebeccapink6398):

OK :D

OpenStudy (31356):

Other questions I meant

OpenStudy (rebeccapink6398):

wait I have another question

OpenStudy (31356):

Sure :D

OpenStudy (rebeccapink6398):

How much grapefruit juice is needed to make 32 servings of fruit punch? A. 12 c B. 9 c C. 8 c D. 6

OpenStudy (rebeccapink6398):

I don't get it :(

OpenStudy (31356):

A) 12 cups :D

OpenStudy (31356):

Just get one of the numbers then Set up a ratio of 3/4=?/16

OpenStudy (rebeccapink6398):

:D

OpenStudy (31356):

16*3 then divide by 4

OpenStudy (rebeccapink6398):

THX! :D :D :D triple s=super smilies

OpenStudy (rebeccapink6398):

oops! lol typos

OpenStudy (31356):

Yea, lol Triple smileys :D:D:D

OpenStudy (anonymous):

on the second of you really just need to times it by 4 8 * 4 = 32 4 * 4=16 3 * 4 = 12 1 * 4 =4